OpenStudy (anonymous):

find the value of tan 1/2θ when tanθ is 3/4

OpenStudy (anonymous):

so the first step is to take the inverse tan of 3/4; this will give u the angle u need. Then sub that answer into 1/(2*phi) and solve. Give it a try

OpenStudy (hba):

@Nutty First of all you have to find the value of θ You are given that, \[\tanθ =3/4\] We have to find, \[\tan \frac{ 1 }{ 2\theta }=?\] I would suggest yoiu to start out with finding θ. @Nutty I hope it is clear ?

OpenStudy (anonymous):

im guessing the question is asking for \[\tan (\frac{ 1 }{ 2 }\theta)\] so he'll need to use half angle formulae for sin and cos. and from tan(theta) = 3/4 you should be able to tell what sin(theta) and cos(theta) are.

OpenStudy (raden):

use the formula :